# Signs as Tokens: An Autoregressive Multilingual Sign Language Generator

### v1

@@ Abstract
Sign language is a visual language that encompasses all linguistic features of natural languages and serves as the primary communication method for the deaf and hard-of-hearing communities. While many studies have successfully adapted pretrained language models (LMs) for sign language translation (sign-to-text), drawing inspiration from its linguistic characteristics, the reverse task of sign language generation (SLG, text-to-sign) remains largely unexplored. Most existing approaches treat SLG as a visual content generation task, employing techniques such as diffusion models to produce sign videos, 2D keypoints, or 3D avatars based on text inputs, overlooking the linguistic properties of sign languages. In this work, we introduce a multilingual sign language model, Signs as Tokens (SOKE), which can generate 3D sign avatars autoregressively from text inputs using a pretrained LM. To align sign language with the LM, we develop a decoupled tokenizer that discretizes continuous signs into token sequences representing various body parts. These sign tokens are integrated into the raw text vocabulary of the LM, allowing for supervised fine-tuning on sign language datasets. To facilitate multilingual SLG research, we further curate a large-scale Chinese sign language dataset, CSL-Daily, with high-quality 3D pose annotations. Extensive qualitative and quantitative evaluations demonstrate the effectiveness of SOKE. The project page is available at https://2000zrl.github.io/soke/.
Figure 1: In this paper, we propose Signs as Tokens (SOKE), a unified sign language generator that can produce multilingual signs from text inputs. (Left: Chinese sign language; Right: American sign language.)
@@ 1 Introduction
Sign language is the primary form of communication method utilized by the deaf and hard-of-hearing communities, embodying all linguistic properties found in spoken languages, including discrete semantic units and grammatical structure [68, 13]. However, despite its importance, communication between hearing and hard-of-hearing communities still relies heavily on trained interpreters or proficient individuals which often exacerbates their social exclusion, creating barriers in everyday interactions. As an attempt to mitigate the barrier between sign and spoken languages, several methods have been developed to tackle sign language processing by unfolding the task into two research problems: sign language translation (SLT, sign-to-text) [6, 84, 24, 77, 64] and sign language generation (SLG, text-to-sign) [50, 3, 69, 83, 57]. Numerous research efforts have established significant progress in SLT by leveraging 2D and 3D representations coupled with large-scale pretrained language models (LMs) [6, 13, 49, 24, 64, 71, 17].
Despite the advent of SLT methods, the task of SLG remains relatively unexplored. Current methods approach SLG as a visual content generation task, leveraging GANs [57, 54] or diffusion models [3, 43] thereby neglecting the linguistic nature of sign languages, leading to suboptimal performance and missing benefits of pretrained LMs such as generalizability and scalability [60, 26, 16]. In particular, sign languages share the fundamental linguistic properties and the discrete structure of natural spoken languages [68, 13], which necessitates processing using language modeling techniques, rather than treating them as a continuous motion generation task. A common approach to discretize sign languages is employing glosses–the written form of signs–as an intermediate representation [83, 54, 20, 59]. While glosses provide a step towards more natural modeling of sign languages, they require meticulous annotation efforts and impose a predefined information bottleneck that may fall short of fully capturing sign semantics. In this work, we approach SLG using a discrete sign space to effectively capture the linguistic structure of sign languages. To do so, we build a sign tokenizer that learns a mapping between continuous motions and discretized tokens using a compact quantized latent space. With such formulation, we can frame the generative task in an autoregressive manner, akin to how natural languages are processed.
In this work, we introduce SOKE, a multilingual sign language generation model designed to produce 3D signs in an autoregressive manner. Unlike body motion modeling, sign languages utilize both manual and non-manual parameters to convey information [29, 78, 81, 23]. Previous works [69, 67, 3] usually overlook the multi-cue nature of sign languages, treating SLG as a whole-body 3D motion generation task. In contrast, we propose a decoupled sign tokenizer that models hand articulations and body posture using distinct tokens. This approach enhances 3D sign modeling and effectively improves the generalization of the sign-token mapping. Given the limited vocabulary size in continuous sign language datasets, we leverage a large-scale pretrained multilingual LM [32] to enhance generalization to an unconstrained domain of discourse. Additionally, to support multilingual SLG in a unified way, we fine-tune the proposed model on a multilingual sign language dataset that includes American and Chinese sign languages by incorporating a series of special tokens to specify the target sign language and body part. To sum up, the contributions of this paper can be summarized as:
@@
  • We propose Signs as Tokens (SOKE), a novel approach for sign language generation. With the aid of a pretrained LM, SOKE can deal with multiple sign languages in a unified model.
  • Motivated by the multi-cue property of sign languages, we propose a decoupled sign motion tokenizer (DETO) that can effectively capture semantics conveyed by various body parts.
  • We curate a large-scale Chinese sign language dataset, CSL-Daily [79], with accurate SMPL-X [40] pose annotations.
@@
  • We demonstrate that SOKE achieves state-of-the-art performance on two challenging benchmarks, How2Sign [8] and CSL-Daily [79], using a single unified model.
Figure 2: An overview of our proposed method, Signs as Tokens (SOKE). Our approach consists of two key components: a decoupled tokenizer (DETO) and an autoregressive multilingual generator (AMG). DETO transforms the input pose sequence to a set of discrete tokens using three distinct vector quantized variational autoencoders, each one composed of a motion encoder ($\mathcal{E}$), a motion decoder ($\mathcal{D}$), and a codebook. These components work together to discretize input sign sequences from various body parts. Leveraging a pretrained language model (LM), the AMG can later generate multilingual signs from text prompts with a tailored part-wise decoding strategy.
@@ 2 Related Work
Sign Language Processing. There are various research topics in sign language processing, including sign language recognition (SLR) [36, 27, 18, 17, 82, 12], sign language translation (SLT) [5, 28, 6, 66, 13, 72], and sign language generation (SLG, also known as sign language production) [9, 53, 54, 83, 69, 2]. Among these, SLT and SLG are two key tasks that focus on sign-to-text and text-to-sign conversion, respectively, constituting a two-way communication system between the deaf and hearing communities. Motivated by the linguistic nature of sign languages, a significant number of works [6, 64, 24] have integrated LMs pretrained on large natural language corpora into SLT pipelines and observed significant performance improvements.
@@
In contrast with SLT, SLG remains largely unexplored. Gloss-based methods [83, 54, 70, 65, 19, 59, 20] approach SLG by utilizing glosses as intermediate representations, following a text-to-gloss-to-sign pipeline. However, glosses act as an information bottleneck, which limits the generation performance. Several works [2, 50, 52] propose to directly regress 2D joints to represent sign poses, simplifying the modeling process, however, they suffer from mode collapse due to the limited scale of data. Recent state-of-the-art SLG works can be categorized into two classes: the first group of methods [3, 9] employ diffusion models to generate sign motions conditioned on text inputs; the second group of methods [7, 69, 67] consider the linguistic nature of sign languages and adopts a tokenizer-LM two-stage autoregressive generation approach. Nevertheless, few works consider the multi-cue property of sign languages which limits their expressivity. In addition, there has been a notable gap in the literature on unified models for text-driven multilingual SLG. In this work, we propose a decoupled tokenizer that maps the body and both hands into separated quantized latent spaces and address multilingual SLG in a unified model with pure text inputs.
Sign Tokenizers. Besides using glosses as discrete sign representations [83, 70], several studies [69, 67, 13, 35] focus on training sign tokenizers using purely visual data. For instance, SignVAE [69] employs a basic VQ-VAE [61, 15] to tokenize sign motions, while DVQ-VAE [67] introduces a dynamic quantizing technique to adjust encoding length. In the field of SLT, VQ-Sign [13] tokenizes sign videos at both the character and word levels. In contrast to the previous methods, we propose a simple yet effective decoupled tokenizer, which is deeply embedded with the task of SLG, considering the multi-cue nature of sign languages.
Motion Generation. Motion generation is an important area of research with substantial impacts on domains such as human-computer interaction and robotics. Early studies [14, 41] focus on generating motions based on predefined action categories, which constrain their practical applications. With the emergence of vision-language models [44], recent motion generation methods have shifted to using text as a condition, making them more accessible and intuitive. These methods can be categorized into two main types: 1) Diffusion-based approaches [75, 56, 4], which utilize latent diffusion [47] to generate motions from Gaussian noises, incorporating text conditions through a cross-attention mechanism. 2) Autoregressive approaches [22, 76, 80, 74, 15], which draw inspiration from text generation methods by first tokenizing motion sequences and then using language models to perform autoregressive generation.
Applying motion generation methods in the domain of SLG has demonstrated particularly promising results [3, 67]. However, a major limitation of these studies is their insufficient attention to hand articulations. In this work, we propose a sign-language-specific tokenizer to emphasize on complex hand articulations.
@@ 3.1 Data Preparation
Although NSA [3] and SignAvatars [69] have made efforts in curating high-quality (pseudo) SMPL-X annotations for an American Sign Language (ASL) dataset, How2Sign [8], there is a notable lack of 3D annotations for other continuous sign language datasets, impeding multilingual SLG research. In this work, we rely on a two-stage approach, leveraging state-of-the-art 3D pose estimation methods [31, 42], to extract high-fidelity SMPL-X poses for CSL-Daily [79], the largest Chinese Sign Language (CSL) dataset. Specifically, we initially estimate a set of rough 3D body poses for each signer using OSX [31] along with accurate 3D hand poses from WiLoR [42]. Given that OSX often falls short in accurately capturing the arm poses, we follow [3] and refine the upper body joint rotations by minimizing a re-projection loss between the estimated joints and the detected 2D joint keypoints from Mediapipe [34]. Using this process we can ensure accurate whole-body annotations. We represent each sign motion sequence as $\mathbf{S}\in\mathbb{R}^{T\times d}$, where $T$ is the sequence length and $d=133$ denotes the number of SMPL-X parameters, including 11 upper body joints, 30 hand joints, and 10 expression parameters. For additional details, we refer the reader to the supplementary materials.
3.2 Framework Overview
An overview of our method is illustrated in Figure 2. The method is comprised of two main components: a decoupled tokenizer (DETO) and an autoregressive multilingual generator. DETO includes three vector quantized variational autoencoders (VQ-VAEs) [61, 22], each corresponding to one of three informative regions: the upper body ($B$, which includes the face by default), the left hand ($LH$), and the right hand ($RH$). Given a sign motion sequence with $T$ frames, denoted as $\mathbf{S}=\{s_{i}\}_{i=1}^{T}$, where $s_{i}\in\mathbb{R}^{d}$, DETO initially encodes $\mathbf{S}$ into three token sequences: $\mathbf{\hat{Z}}^{B}_{1:N}$, $\mathbf{\hat{Z}}^{LH}_{1:N}$, and $\mathbf{\hat{Z}}^{RH}_{1:N}$, where $N=T/\alpha$ and $\alpha$ represents a temporal downsampling factor. These encoded features are then fed into decoders to obtain the reconstructed motions $\mathbf{\hat{S}}$.
Once DETO is trained, we convert all sign motion sequences into their corresponding token representations. The derived tokens are collected and integrated into the raw text vocabulary of a pretrained multilingual language model (LM), mBART [32], for subsequent supervised fine-tuning on sign language datasets [8, 79]. Finally, during inference, given a text prompt $\mathbf{X}=\{x_{i}\}_{i=1}^{L}$ from any of the supported languages, the LM can generate sign motion tokens. The full 3D avatar motions can be generated by feeding the derived tokens into the DETO decoders.
@@ 3.3 Decoupled Tokenizer
Figure 3: Workflow of our decoupled tokenizer (DETO). It is composed of three parallel VQ-VAEs, each dedicated to generating motion tokens for a different part of the signer’s body: the upper body, left hand, and right hand.
To discretize continuous signs into tokens, we develop a sign tokenizer based on the well-established VQ-VAE [61, 22]. Existing motion generation research primarily focuses on body movements, such as running and jumping, while often neglecting the role of hands. However, in sign language, hands are crucial for conveying information [17, 55]. To bridge this gap, we propose DETO, a decoupled tokenizer that utilizes three VQ-VAEs to simultaneously model key regions: the upper body and both hands. By assigning two of the VQ-VAEs specifically to the hands, we can independently model the hand motions and effectively improve the expressivity of the model. An illustration of DETO is shown in Figure 3.
Given a sign motion input $\mathbf{S}\in\mathbb{R}^{T\times d}$, we first decompose it into three part-wise motion sequences based on kinematic tree of SMPL-X [40]: $\mathbf{S}^{p}\in\mathbb{R}^{T\times d_{p}}$, where $p\in\{B,LH,RH\}$. Moreover, we build three distinct VQ-VAEs, where each of them consists of an encoder $\mathcal{E}_{p}(\cdot)$, a decoder $\mathcal{D}_{p}(\cdot)$, and a learnable codebook $\mathbf{Z}^{p}\in\mathbb{R}^{N_{Z}^{p}\times C}$, where $N_{Z}^{p}$ represents the number of codes and $C$ denotes the code dimension. For each motion sequence, the corresponding encoder first projects the sequence into a latent space: $\mathbf{S}_{f}^{p}=\mathcal{E}_{p}(\mathbf{S}^{p})\in\mathbb{R}^{N\times C}$, using a stack of 1D-CNN layers. Then, for each pose we can derive a set of discrete tokens $\mathbf{\hat{Z}}^{p}=\{\hat{z}_{i}^{p}\}_{i=1}^{N}$ using a quantizer $Q(\cdot)$, which searches the nearest neighbor from the codebook $\mathbf{Z}^{p}$:
@@
We then feed the obtained token sequence to the corresponding part-decoder $\mathcal{D}_{p}$ to reconstruct the input motions: $\mathbf{\hat{S}}^{p}=\mathcal{D}_{p}(\mathbf{Z}^{p})$. Following [22], the objective function of each VQ-VAE is defined as: $\mathcal{L}_{vq}^{p}=\mathcal{L}_{rec}^{p}+\mathcal{L}_{emb}^{p}+\mathcal{L}_{com}^{p}$, where $\mathcal{L}_{rec}^{p}=\|\mathbf{\hat{S}}^{p}-\mathbf{S}^{p}\|_{2}^{2}$ denotes the reconstruction loss, $\mathcal{L}_{emb}^{p}$ is the embedding loss, and $\mathcal{L}_{com}^{p}$ is the commitment loss, respectively. The overall loss function for the entire DETO is defined as the sum of $\mathcal{L}_{vq}^{p}$ across all body parts.

### v3

# Signs as Tokens: A Retrieval-Enhanced Multilingual Sign Language Generator

@@ Abstract
Sign language is a visual language that encompasses all linguistic features of natural languages and serves as the primary communication method for the deaf and hard-of-hearing communities. Although many studies have successfully adapted pretrained language models (LMs) for sign language translation (sign-to-text), the reverse task—sign language generation (text-to-sign)—remains largely unexplored. In this work, we introduce a multilingual sign language model, Signs as Tokens (SOKE), which can generate 3D sign avatars autoregressively from text inputs using a pretrained LM. To align sign language with the LM, we leverage a decoupled tokenizer that discretizes continuous signs into token sequences representing various body parts. During decoding, unlike existing approaches that flatten all part-wise tokens into a single sequence and predict one token at a time, we propose a multi-head decoding method capable of predicting multiple tokens simultaneously. This approach improves inference efficiency while maintaining effective information fusion across different body parts. To further ease the generation process, we propose a retrieval-enhanced SLG approach, which incorporates external sign dictionaries to provide accurate word-level signs as auxiliary conditions, significantly improving the precision of generated signs. Extensive qualitative and quantitative evaluations demonstrate the effectiveness of SOKE.
Figure 1: We propose Signs as Tokens (SOKE), a unified sign language generator that can produce multilingual signs from text inputs. (Left: American sign language; Middle: Chinese sign language; Right: German sign language.)
@@ 1 Introduction
Sign language is the primary communication method utilized by the deaf and hard-of-hearing communities, embodying all linguistic properties found in spoken languages, such as discrete semantic units and grammatical structure [77, 19]. As an attempt to mitigate the barrier between sign and spoken languages, several methods have been developed to tackle sign language processing by unfolding the task into two research problems: sign language translation (SLT, sign-to-text) [23, 75, 87, 70, 90, 73, 38, 62] and sign language generation (SLG, text-to-sign) [53, 3, 78, 95, 59].
In contrast with SLT, the usage of language models (LMs) in SLG remains relatively unexplored. Most existing methods approach SLG as a visual content generation task (e.g., video, keypoint, or motion), using GANs [59, 56] or diffusion models [3, 49, 13, 64, 65]. However, these works neglect the linguistic nature of sign languages, leading to suboptimal performance and missing benefits of pretrained LMs such as generalizability and scalability [66, 33, 21]. In particular, sign languages share the fundamental linguistic properties and the discrete structure of spoken languages [77, 19], which necessitates processing using autoregressive language modeling techniques [76].
To formulate SLG in the framework of language modeling, a crucial step is mapping signs into discrete tokens. A common approach to discretizing sign languages involves using glosses—the written form of signs—as an intermediate representation [95, 56, 25, 14]. However, glosses require extensive annotation efforts and impose a predefined information bottleneck, which fail to fully capture the rich semantics of sign languages [44]. To address these limitations and better capture the multi-cue property of sign languages [92, 29, 88], we leverage a vector quantized-variational auto-encoder (VQ-VAE)-based tokenizer [74, 42, 71] that learns mappings between continuous sign motions and discrete tokens for each body part (upper body, left hand, and right hand). During decoding, existing autoregressive motion generation methods [74, 42] typically flatten tokens from different body parts into a single sequence, resulting in inefficient inference due to tripled decoding steps. To overcome this challenge, we propose a novel multi-head decoding method capable of predicting multiple tokens at a time while effectively fusing information from different body parts in real time. Furthermore, given the limited vocabulary size of continuous sign language datasets, we enhance the model’s generalization by instantiating the LM as a multilingual LM [40], enabling it to handle a broader and more diverse domain of discourse. To support multilingual SLG in a unified framework, we train the proposed model on a curated multilingual sign language dataset encompassing American, Chinese, and German sign languages [11, 89, 4].
Recently, retrieval-augmented generation (RAG) [5, 12, 68] has proven effective in enhancing the interpretability and accuracy of content generated by large language models. Similarly, several SLG methods [95, 56] leverage glosses to retrieve sign dictionaries and directly incorporate retrieved signs into the outputs, improving the interpretability of generated signs. However, beyond the inherent limitations of glosses, these methods often face challenges in handling co-articulations between adjacent signs. This is because dictionary signs are typically recorded at the word level, whereas SLG models operate at the sentence level, leading to potential unnaturalness in the generated sequences. To address these issues, we propose a retrieval-enhanced SLG approach that utilizes motion tokens from retrieved dictionary signs as additional conditions for language modeling. This approach avoids the unnaturalness of directly copying signs into model outputs while still benefiting from the precision of dictionary signs. To sum up, the contributions of this paper can be summarized as:
@@
  • We propose a novel SLG approach, Signs as Tokens (SOKE). With the aid of a pretrained LM, SOKE can handle multiple sign languages in a unified model.
  • We propose a multi-head decoding method that reduces inference latency and efficiently fuses information from different body parts, enabling seamless integration of a decoupled tokenizer with the LM.
  • By leveraging external sign dictionaries, we introduce a retrieval-enhanced SLG approach to improve the precision of generated signs.
@@
  • We curate two sign language datasets [89, 4] with accurate SMPL-X [47] poses to facilitate multilingual SLG research. Experiments demonstrate that our SOKE achieves state-of-the-art performance on three challenging benchmarks, How2Sign [11], CSL-Daily [89], and Phoenix-2014T [4], using a single unified model.
Figure 2: An overview of our proposed method, Signs as Tokens (SOKE). We begin by training a VQ-VAE-based decoupled tokenizer to map continuous sign motions into discrete tokens for various body parts (upper body, left hand, and right hand). These sign motion tokens are then integrated into the text vocabulary of a pretrained language model, which serves as the backbone of our autoregressive multilingual generator (AMG). Given a text input, the AMG first retrieves word-level signs from external dictionaries, appends their motion tokens to the text tokens, and feeds them into the language model encoder. During decoding, our novel multi-head decoding strategy generates motion tokens for all body parts simultaneously at each time step. Finally, the derived motion tokens are used to reconstruct sign avatars.
@@ 2 Related Work
Sign Language Generation. Early gloss-based methods [95, 56, 79, 71, 24, 63, 25] approach SLG under a text-to-gloss-to-sign framework. Several works [2, 53, 55] propose to directly regress 2D joints to represent sign poses, simplifying the modeling process, however, they suffer from mode collapse due to the limited scale of data. Recent state-of-the-art SLG works can be categorized into two classes: the first group of methods [3, 13] employs diffusion models to generate sign motions conditioned on text inputs; the second group of methods [10, 78, 76] adopts a tokenizer-LM two-stage autoregressive generation approach. Among these, [78] incorporates costly gloss and HamNoSys notations [2] to train a semantic adapter, while [10] focuses on generating signs from a single word, limiting their applicability. Moreover, a notable gap exists in the literature regarding unified models for multilingual SLG. In this work, we address these gaps by utilizing a decoupled tokenizer and a pretrained multilingual LM, integrated with our customized multi-head decoding method and retrieval-enhanced approach, to enable text-driven multilingual SLG.
Text-to-Motion Generation. Text-to-motion generation (T2MG) is an important research area with potential applications in human-computer interaction and robotics. Recent T2MG methods can be categorized into two main types: 1) Diffusion-based approaches [82, 57, 6], which utilize latent diffusion [51] to generate motions conditioned on text inputs. 2) Autoregressive approaches [27, 83, 91, 81, 20, 42], which draw inspiration from text generation methods by first tokenizing motion sequences and then using LMs to perform autoregressive generation. When decoding with a decoupled tokenizer [42, 74], these autoregressive approaches simply flatten tokens from various body parts into one sequence, causing addtional decoding steps. In this work, we propose a novel multi-head decoding method that can efficiently predict motion tokens for all body parts at one single step, significantly reducing inference latency.
Retrieval-Enhanced Generation. Using the concept of retrieval to enhance generation models has proven effective across various domains [5, 12, 68, 18, 72, 9]. As a pioneering effort, Piktus et al. [36] propose to retrieve relevant documents before generating answers for knowledge-intensive NLP tasks. FLARE [28] introduces selective retrieval, activating it only when low-confidence tokens appear in the generation process. ToG [61] leverages retrieved knowledge to enhance the reasoning capabilities of large language models. In the field of image generation, retrieval has also been utilized to enable diffusion models to generate out-of-distribution images [58]. In the context of SLG, several methods [95, 56] adopt gloss-based retrieval to obtain dictionary signs and approach SLG in a retrieve-and-concatenate manner. However, glosses have inherent limitations, and such methods often suffer from unrealistic transitions between signs. In this paper, we propose a simple yet effective retrieval-enhanced SLG approach that uses word-level signs as auxiliary conditions, addressing these limitations and improving the overall quality of generations.
@@ 3 Methodology
Figure 3: Comparison of various decoding methods. (a) Sequential decoding flattens all motion tokens into a single sequence, requiring an inefficient $3K$ decoding steps, where $K$ is the target motion sequence length. (b) Parallel decoding improves efficiency by decoding body parts in parallel but lacks information fusion, which may lead to suboptimal performance. (c) Our proposed multi-head decoding balances efficiency and fusion by predicting three tokens simultaneously via different heads, with input embeddings as a weighted average of part-wise token embeddings.
An overview of our method is shown in Figure 2. Our approach mainly consists of two parts: a decoupled tokenizer (DETO) and an autoregressive multilingual generator (AMG). In the first stage, DETO learns mappings between continuous sign motions and discrete tokens for each body part (Section 3.2). In the second stage, AMG is trained to autoregressively generate motion tokens with the proposed multi-head decoding strategy (Section 3.3) conditioned on text inputs and retrieved dictionary signs (Section 3.4).
@@ 3.1 Data Preparation
Although existing SLG works [3, 78] have made efforts in curating high-quality SMPL-X poses for an American sign language (ASL) dataset, How2Sign [11], there is a notable lack of 3D annotations for other continuous sign language datasets, impeding multilingual SLG research. In this work, we rely on a two-stage approach, leveraging state-of-the-art 3D pose estimation methods [39, 48], to extract high-fidelity SMPL-X poses for CSL-Daily [89] and Phoenix-2014T [4], two widely-adopted datasets of Chinese and German sign languages (CSL and DGS), respectively. Specifically, we initially estimate a set of rough 3D body poses for each signer using OSX [39] along with accurate 3D hand poses from WiLoR [48]. Given that OSX often falls short in accurately capturing the arm poses, we follow [3] and refine the upper body joint rotations by minimizing a re-projection loss between the estimated joints and the detected 2D joint keypoints from Mediapipe [43]. Using this process we can ensure accurate whole-body annotations. We represent each sign motion sequence as $\mathbf{S}\in\mathbb{R}^{T\times d}$, where $T$ is the sequence length and $d=133$ denotes the number of SMPL-X parameters, including 11 upper body joints, 30 hand joints, and 10 expression parameters. Please check the supplementary materials for more details and quantitative evaluations of our pose fitting pipeline.
@@ 3.2 Decoupled Tokenizer
Sign languages exhibit a multi-cue property [88, 93, 29, 84, 86, 22], where the semantics of a sign are conveyed simultaneously through body and hand movements. Inspired by this, we leverage a decoupled tokenizer [42, 74, 71], comprising three VQ-VAEs [67], to independently model key regions: the upper body and both hands. An illustration is available in the supplementary materials.
Given a $T$-frame sign motion input $\mathbf{S}\in\mathbb{R}^{T\times d}$, we first decompose it into three part-wise motion sequences based on the kinematic tree of SMPL-X: $\mathbf{S}^{p}\in\mathbb{R}^{T\times d_{p}}$, where $p\in\{B,LH,RH\}$. Moreover, we build three distinct VQ-VAEs, where each of them consists of an encoder $\mathcal{E}_{p}(\cdot)$, a decoder $\mathcal{D}_{p}(\cdot)$, and a learnable codebook $\mathbf{Z}^{p}\in\mathbb{R}^{N_{Z}^{p}\times C}$, where $N_{Z}^{p}$ represents the number of codes and $C$ denotes the code dimension. For each motion sequence, the corresponding encoder first projects the sequence into a latent space: $\mathbf{S}_{f}^{p}=\mathcal{E}_{p}(\mathbf{S}^{p})=\{s_{f,i}^{p}\}_{i=1}^{T_{f}}\in\mathbb{R}^{T_{f}\times C}$, using a stack of 1D-CNN layers. Then, for each pose we can derive a set of discrete tokens $\mathbf{\hat{Z}}^{p}=\{\hat{z}_{i}^{p}\}_{i=1}^{T_{f}}$ using a quantizer $Q(\cdot)$, which searches the nearest neighbor from the codebook $\mathbf{Z}^{p}$:
@@
We then feed the obtained token sequence to the corresponding part-decoder $\mathcal{D}_{p}$ to reconstruct the input motions: $\mathbf{\hat{S}}^{p}=\mathcal{D}_{p}(\mathbf{Z}^{p})$. Following [27], the objective function of each VQ-VAE is defined as: $\mathcal{L}_{vq}^{p}=\mathcal{L}_{rec}^{p}+\mathcal{L}_{emb}^{p}+\mathcal{L}_{com}^{p}$, where $\mathcal{L}_{rec}^{p}=\|\mathbf{\hat{S}}^{p}-\mathbf{S}^{p}\|_{2}^{2}$ denotes the reconstruction loss, $\mathcal{L}_{emb}^{p}$ is the embedding loss, and $\mathcal{L}_{com}^{p}$ is the commitment loss, respectively.